# StrucText-Eval: An Autogenerated Benchmark for Evaluating Large Language Model’s Ability in Structure-Rich Text Understanding

### v2

@@ Abstract
Given the substantial volumes of structured data held by many companies, enabling Large Language Models (LLMs) to directly understand structured text in non-structured forms could significantly enhance their capabilities across various business scenarios. To this end, we propose evaluation data generation method for assessing LLM’s ability in understanding the structure-rich text, which generates structured data of controllable complexity based on manually crafted question templates and generation rules. Building on this generation method, we introduce StrucText-Eval, a benchmark comprising 6,032 questions across 8 different structured languages and 29 specific tasks. Furthermore, considering human proficiency in rule-based tasks, we also present StrucText-Eval-Hard, which includes 3,016 questions designed to further examine the gap between LLMs and human performance. Results indicate that the best-performing LLM currently achieve an accuracy of 65.0% on StrucText-Eval-Hard, while human accuracy reaches up to 95.7%. Moreover, while fine-tuning using StrucText-Eval can enhance existing LLMs’ understanding of all structured languages, it does not necessarily improve performance across all task types. The benchmark and generation codes are open sourced in https://github.com/MikeGu721/StrucText-Eval
@@ 1 Introduction
Structured data, often represented by various structured languages such as JSON [19], YAML [8], ORG [11], or Markdown [13], Latex [15] etc., has consistently been central to corporate data strategies due to its ability to capture, store, and analyze essential information systematically. The inherent benefits of structured data lie in its standardized format and high degree of organization, which facilitates efficient data querying and machine processing, clearly surpassing the inherent chaos of unstructured data. However, with the advancement of large language models (LLMs) [1, 26, 27, 22], there has been a significant shift towards the effective utilization of unstructured data, attributed to the LLMs’ capacity to comprehend and generate complex and nuanced semantics within such data [3]. Considering that structured data can be directly presented in an unstructured format, it makes us wonder: whether it is possible to rely on LLMs to interpret structured data through unstructured format directly.
@@
Current LLMs researches have addressed their comprehension of structure-rich text, with a primary focus on the understanding of limited categories: Graphs [9, 18, 14, 24, 6], Tables [21, 4, 17] and JSON [5, 23]. However, these categories do not encompass all potential use cases of structure-rich text. For instance, scenarios requiring direct understanding of articles in Latex or Markdown formats, data in YAML or ORG formats, or various custom structured languages are not adequately covered. Moreover, existing benchmarks often rely on manually annotated data for evaluation, which limits the development of robust evaluation frameworks and potentially facilitates model cheating [31].
@@
In response, we propose a method for automatically generating evaluation data to assess models’ capabilities in understanding structure-rich text. This method is applied to 8 structured languages, as shown in Fig. 1, across 29 specific tasks, enabling data generation with controllable difficulty by adjusting the depth of structured nesting. Based on this method, we further introduce the Structure-Rich Text Evaluation Benchmark (StrucText-Eval), a comprehensive benchmark with 6,032 samples designed to evaluate the proficiency of LLMs in deciphering embedded structures within input text. StrucText-Eval aims to determine whether LLMs understand raw structural tags, execute logical inferences based on the decoded semantics of these symbols, and construct their responses according to instruction requirements. Considering that humans excel at understanding structured expressions, we further propose StrucText-Eval-Hard, a subset of StrucText-Eval consisting of 3,012 samples with the longest questions, highlighting the disparity between LLMs and human understanding of structured data.
The experiments show that StrucText-Eval significantly challenges current LLMs. The highest accuracy achieved by LLMs using various prompt methods was 78.4%, dropping to 65.0% on StrucText-Eval-Hard. In contrast, human participants reached 95.7% accuracy on StrucText-Eval-Hard, underscoring the inadequacy of existing LLMs in processing complex structural information. Utilizing StrucText-Eval data enhances the capabilities of current closed-source LLMs and improves open-source LLMs by fine-tuning them with this benchmark. This finding confirms StrucText-Eval’s critical role in evaluating and optimizing model performance. However, more than using the training data from StrucText-Eval alone is required to surpass GPT4-Turbo’s performance on this test suite. Additionally, not all tasks can be improved solely by training on the relevant dataset. Therefore, StrucText-Eval provides a comprehensive evaluation framework for assessing LLMs’ ability to understand structured text, leaving room for further improvement.
We present the following contributions to our research: (1) We developed a method for generating evaluation data for eight structured languages. This method allows control over the complexity of the generated evaluation data by manipulating the quantity of structural nesting. (2) We introduce StrucText-Eval, a comprehensive benchmark comprising 6,032 samples across eight structured languages and 29 evaluation tasks. This benchmark is designed to assess LLMs’ capabilities in understanding structured data comprehensively. (3) We have conducted extensive experiments, including evaluations of existing large models’ capabilities in understanding structure-rich text and enhancements to the capacity of open-source LLMs in this area. Our findings indicate a significant gap between current LLMs and humans in understanding structure-rich text.
@@ 2 Related Work
2.1 Structural Text Understanding Enhanced Models
Previous explorations over structural information focused on tasks related to programming languages and graphs. Mou et al. [16] built a tree-based convolutional neural network that outperformed language models on functionality classification tasks and pattern detection tasks over programming languages. Wang et al. [29] outperformed state of the art (SOTA) on code clone detection benchmarks, applying graph neural networks on abstract syntax trees (ASTs) representation augmented by control and data flow edges, which is different in architecture from language models curated for natural language tasks. Zhou et al. [32] built a graph neural network over rich node representations of code and gained SOTA on vulnerability identification in 4 large-scaled open-source C projects. Perozzi et al. [18] proposed a dedicated graph neural network (GNN) to encode graph into embedding space of LLMs, and used LLMs as an decoder to solve the graph tasks in GraphQA [9]. Chen et al. [6] explored LLMs’ potential in learning graphs, focusing on node classification tasks, using LLMs as enhancers for GNN predictor and directly as predictor. Chen et al. [5] disclosed the employment of formats other than natural language in CoT reasoning and multi-agent communication significantly enhances model performance while also decreasing token usage on reasoning tasks.
2.2 Structural Text Understanding Evaluation Datasets
The similar datasets dedicating to structure information understanding are scarce. GraphQA [9] is a benchmark for assessing LLMs on reasoning over graph data, composed of 7 simple graph tasks, with exploration on methods in encoding graph-structured data as text. Struc-Bench [25] collect data from open source to evaluate if LLM can generate complex structured tabular text. Sui et al. [21] consists of 8 tasks on table-formatted input texts, i.e., input data organized in row-column structures, with studying in input design and prompt enhancement. Pasupat and Liang [17] exposed a dataset of 22,033 complex questions on Wikipedia tables and proposed a parsing algorithm for answering complex questions on semi-structured tables. For tabular data, Shwartz-Ziv and Armon [20] showed that deep models were not a good solution; Gorishniy et al. [12] designed two specialized baseline architecture for tabular data and concluded that universally superior solution does not exist. Wang et al. [28] is another graph-based problem solving benchmark, with graph structure described with natural language. Fatemi et al. [9] and Sui et al. [21] include comparison of encoding methods for structure information, i.e., graphs and tables, respectively, showing that LLMs performance varies on different encodings and the relatively vacancy on LLM’s understanding of structure-rich texts input.
Our dataset diverges from previous work in two significant ways. Firstly, our tasks exclusively concentrate on analyzing structural information and executing dependent inference, unlike prior approaches [16, 32] that employ structural data primarily as an auxiliary tool for classification or semantic comprehension. StrucText-Eval focus on symbolic structure by erasing content semantics from input texts, preserving only structural details, which renders the texts semantically obscure and the tasks semantically agnostic. Secondly, our benchmark extends beyond traditional graph reasoning and table information retrieval to include a broader spectrum of structure-rich text types. It also incorporates comprehensive structure-based inference tasks that are tailored to specific input formats.
@@ 3.2 Dataset Generation
An example of JSON’s PathCompose is shown in Fig. 2 to illustrate the dataset generation process. The generation process mainly entails constructing an abstract structure tree, manually drafting question templates, and developing corresponding answer discovery algorithms. The primary objective of constructing the abstract structure tree is establishing a hierarchically nested structured data template. The complexity of the sample increases when more nodes are set in this tree.
Eight task categories have been delineated for eight languages, as detailed in Fig. 0(a). 29 rules and question templates have been formulated for these tasks, with the specific rule templates detailed in Sec. F in the Appendix. Each sample in the dataset comprises four main fields: “Reference”, “Question”, “Requirement” and “Answer”. We give examples for each language and task in Sec. E in the Appendix.
@@ 3.3 Statistic Information
StrucText-Eval has assembled a dataset comprising 6,032 samples, with each of the 29 specific tasks for eight languages containing 208 samples. These tasks are divided into “simple” and “hard” categories based on data length and complexity, as depicted in Fig. 3(a). The dataset is partitioned into training and testing sets at a 9:2 ratio, which supports extensive training while also enabling the evaluation of model generalization on significant volumes of unseen data. Detailed statistics regarding the number of samples, lengths, and complexity levels across all tasks, languages and difficulties are detailed in Tab. 3 in Sec. C in Appendix.
@@ 4.1 Models
We tested five Close-Source LLMs accessed through public APIs: GPT4, GPT4 Turbo, MINIMAX, SPARK, and ERNIE. Besides these API-Based models, we also finetuned Llama-7B-v1 under several conditions with partially frozen parameters. The API calling and finetuning setting is listed in Appendix C.
@@ 4.2 Prompt-based Method
We assessed the influence of prompt design on LLM performance by employing four distinct prompt settings, and the detailed implementation of these prompts can be found in Tab. LABEL:tab:prompt in the Appendix. Naive: This setting involves appending the question directly after the input text, serving as a straightforward task formulation. CoT: In this configuration, the LLM generates its thinking process before directly response to the question, aimed at facilitating task resolution [30]. Few-Shot Demonstration: This approach leverages the principles of few-shot learning [10], with variations including one and two example instances to guide the model. Rule Hint: This setup incorporates manually crafted, step-by-step guidance appended after the input text to aid the inference process.
4.3 Finetune-based Method
We further explored the efficacy of finetuning OpenLLaMA 7B v1 using various subsets of data. Our finetuning configurations are described as follows: Base: Utilizes the core StrucText-Eval dataset, referred to as the base StrucText-Eval setting. 2-Shot: Implements a few-shot (2-shot) learning strategy within the StrucText-Eval framework. Rule Hint: Adapts StrucText-Eval by appending a new task-specific rule hint to each sample, effectively redefining the input text as a combination of the question, answer, and original input, with the answer delineated as a step-by-step procedure. Mix: is a variant of Base with half of samples replaced by samples from ShareGPT111https://huggingface.co/datasets/shibing624/sharegpt_gpt4.
@@ 4.4 Evaluation Method
This paper employs three distinct evaluation methodologies: GPT-Judge and RougeL. GPT-Judge utilizes GPT4-turbo to assess whether the LLM output semantically aligns with the ground truth. RougeL measures the degree of character similarity between them. Given that the RougeL expects LLM to output no superfluous content, it imposes additional demands on the LLM’s ability to adhere to instructions, potentially distorting the evaluation results. Therefore, we primarily present the results of GPT-Judge in the main text, while the outcomes of RougeL are listed in Sec. G in the Appendix.
@@ 5 Analysis
Ernie-turbo | Base | 27.4 | 0.2 | 0.0 | 0.0 | 13.0 | 2.6 | 14.3 | 24.4 | 2.8 | 22.3 | 23.2 | 11.3 | 0.5 | 11.1 | 5.4 | 1.8 | 12.5 |
CoT | 19.4 | 0.1 | 0.1 | 0.2 | 10.9 | 1.5 | 13.4 | 19.7 | 0.8 | 6.7 | 26.6 | 6.2 | 1.4 | 6.2 | 1.8 | 0.0 | 7.2 | |
1-shot | 25.0 | 0.0 | 4.2 | 8.7 | 21.9 | 12.5 | 8.3 | 25.0 | 12.5 | 20.0 | 8.3 | 19.4 | 0.0 | 0.0 | 25.0 | 0.0 | 15.2 | |
2-shot | 12.5 | 12.5 | 12.5 | 8.3 | 21.9 | 20.8 | 16.7 | 5.0 | 4.2 | 10.0 | 4.2 | 18.3 | 25.0 | 0.0 | 25.0 | 25.0 | 13.4 | |
Rule Hint | 32.5 | 0.0 | 0.0 | 0.0 | 9.4 | 0.0 | 12.5 | 32.5 | 8.3 | 20.0 | 25.0 | 14.4 | 0.0 | 6.2 | 0.0 | 0.0 | 13.8 | |

### v3

# StrucText-Eval: Evaluating Large Language Model’s Reasoning Ability in Structure-Rich Text

@@ Abstract
The effective utilization of structured data, integral to corporate data strategies, has been challenged by the rise of large language models (LLMs) capable of processing unstructured information. This shift prompts the question: can LLMs interpret structured data directly in its unstructured form? We propose an automatic evaluation data generation method for assessing LLMs’ reasoning capabilities on structure-rich text to explore this. Our approach supports 8 structured languages and 29 tasks, generating data with adjustable complexity through controllable nesting and structural width. We introduce StrucText-Eval, a benchmark containing 5,800 pre-generated and annotated samples designed to evaluate how well LLMs understand and reason through structured text. StrucText-Eval is divided into two suites: a regular Test suite (3,712 samples) and a Test-Hard suite (2,088 samples), the latter emphasizing the gap between human and model performance on more complex tasks. Experimental results show that while open-source LLMs achieve a maximum accuracy of 74.9% on the standard dataset, their performance drops significantly to 45.8% on the harder dataset. In contrast, human participants reach an accuracy of 92.6% on StrucText-Eval-Hard, highlighting LLMs’ current limitations in handling intricate structural information. The benchmark and generation codes are open sourced in https://github.com/MikeGu721/StrucText-Eval
StrucText-Eval: Evaluating Large Language Model’s Reasoning Ability in Structure-Rich Text
Zhouhong Gu♡††thanks: Equal Contribution , Haoning Ye♡∗, Xingzhou Chen♡, Zeyang Zhou♡, Hongwei Feng♡††thanks: Corresponding authors. , Yanghua Xiao♡♣†, ♡Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University ♣Fudan-Aishu Cognitive Intelligence Joint Research Center {zhgu22, xzchen24}@m.fudan.edu.cn {hnye19,zeyangzhou20,hwfeng,shawyh}@fudan.edu.cn
@@ 1 Introduction
Structured data, often represented by various structured languages such as JSON Pezoa et al. (2016), YAML Evans (2001), ORG org (2023), or Markdown Gruber (2012), Latex Lamport (1985) etc., has consistently been central to corporate data strategies due to its ability to capture, store, and analyze essential information systematically. The inherent benefits of structured data lie in its standardized format and high degree of organization, which facilitates efficient data querying and machine processing, clearly surpassing the inherent chaos of unstructured data. However, with the advancement of large language models (LLMs) Achiam et al. (2023); Touvron et al. (2023a, b); Sun et al. (2021), there has been a significant shift towards the effective utilization of unstructured data, attributed to the LLMs’ capacity to comprehend and generate complex and nuanced semantics within such data Brown et al. (2020). Considering that structured data can be directly presented in an unstructured format, it makes us wonder: whether it is possible to rely on LLMs to interpret structured data through unstructured format directly.
@@
Current LLM researchers have addressed their comprehension of structure-rich text of limited categories: Graphs Fatemi et al. (2023); Perozzi et al. (2024); Guo et al. (2023); Tang et al. (2023a); Chen et al. (2023), Tables Sui et al. (2024); Campbell-Kelly (2003); Pasupat and Liang (2015) and JSON Chen et al. (2024); Suzgun et al. (2022). However, these categories do not encompass all potential use cases of structure-rich text. For instance, scenarios requiring a direct understanding of articles in Latex or Markdown formats, data in YAML or ORG formats, or various custom-structured languages need to be adequately covered. Moreover, existing benchmarks often rely on manually annotated data for evaluation, which limits the development of robust evaluation frameworks and potentially facilitates model cheating Zhou et al. (2023).
@@
We propose a method for automatically generating evaluation data to assess models’ capabilities in structure-rich text reasoning. This method is applied to 8 structured languages, as shown in Fig. 1, across 29 specific tasks, enabling data generation with controllable difficulty by adjusting the depth of structured nesting and the number of width and columns in the sample. Based on this method, we further introduce the Structure-Rich Text Evaluation Benchmark (StrucText-Eval), a comprehensive benchmark with 5,800 pre-generated and annotated samples designed to evaluate the proficiency of LLMs in deciphering embedded structures within input text. StrucText-Eval aims to evaluate whether LLMs understand raw structural tags, execute logical inferences based on the decoded semantics of these symbols, and organize their responses according to instruction requirements. StrucText-Eval is divided into Test suite and Test-Hard suite, each consisting of 3,712 and 2,088 samples. Considering that humans excel at understanding structured expressions, The Test-Hard suite features significantly longer questions, with an average length of 16,535 characters, and longest category contains 102,531 characters, which underscores the considerable gap between LLMs and human capabilities in comprehending and processing structured data. The experimental results indicate that StrucText-Eval presents significant challenges in evaluating current LLMs’ structured text processing capabilities. While various open-sourced models achieve a maximum accuracy of 74.9% under different prompting methods, their performance declines markedly to 45.8% when tested on the more complex StrucText-Eval-Hard dataset. In contrast, human participants attain an accuracy of 92.6% on StrucText-Eval-Hard, highlighting the limitations of existing LLMs in comprehending and reasoning through complex structural information.
@@ 2 Related Work
[…]
2.1 Structural Text Understanding Enhancements
Recent efforts to enhance LLMs have focused on integrating external structures such as graphs, tool flows, and cross-domain representations to improve reasoning capabilities across various tasks. For instance, ControlLLM utilizes tool graphs to decompose complex multimodal tasks, resulting in enhanced performance on image and audio processing tasks by leveraging the topological dependencies of tools Liu et al. (2023). Graph-based models like GraphGPT and BooG have shown promising results, with the former improving generalization across node classification and molecular tasks via graph instruction tuning Zhao et al. (2023); Tang et al. (2024). At the same time, the latter employs virtual supernodes to unify graph structures across domains, fostering cross-domain task transferability Cheng et al. (2024). Additionally, methods like RC2R demonstrate the effective combination of knowledge graphs and LLMs for domain-specific causal reasoning, particularly in financial risk propagation tasks Yu et al. (2024). These advancements highlight the benefits of embedding structural elements, from graph architectures to domain-specific knowledge graphs, within LLM frameworks to improve task-specific inference and reasoning.
2.2 Structural Text Understanding Evaluation
Evaluating LLMs’ understanding of structured data has become increasingly critical, though benchmarks remain limited. GraphQA and Struc-Bench are key datasets that assess LLMs’ reasoning over graph-structured data and tabular text, respectively, illustrating the models’ varying capabilities based on input encoding Fatemi et al. (2023); Tang et al. (2023b). More specialized benchmarks, such as TEMPTABQA, evaluate temporal reasoning in tabular data, while TableLLM tests LLMs’ proficiency in handling complex document-based table manipulation tasks Gupta et al. (2023); Zhang et al. (2024). Other works, such as the evaluation of knowledge graph-based reasoning in complex time-series QA systems (JMFRN) Huang et al. (2024), and privacy-oriented graph tasks in GHRatio Yuan et al. (2024), further explore how LLMs handle intricate, structure-rich information, shedding light on their performance across different structured data formats.
Our work diverges from prior research by focusing exclusively on structure-based inference, deliberately removing semantic content to challenge LLMs to reason purely from structural patterns. Unlike previous approaches that use structural data as supplementary input for classification or semantic tasks Pasupat and Liang (2015); Sui et al. (2024), we design semantically agnostic tasks requiring models to infer meaning solely from symbolic structures. Moreover, while earlier benchmarks emphasize graph reasoning or tabular information retrieval, our work extends to a broader spectrum of structure-rich text types, encompassing various input formats and more complex dependency-based inference tasks.
@@ 3.1 Structure-Rich Texts Taxonomy
#Sample | #Reference | #GroundTruth | Depth | Width |
StrucText-Eval-Test | ||||
3,712 | 804 | 47 | - | - |
1,856 | 582 | 19 | 1 | 1 |
1,856 | 1,026 | 74 | 2 | 1 |
StrucText-Eval-Test-Hard | ||||
2,088 | 16,535 | 1,169 | - | - |
232 | 573 | 22 | 1 | 1 |
232 | 614 | 26 | 1 | 2 |
232 | 663 | 25 | 1 | 3 |
232 | 992 | 80 | 2 | 1 |
232 | 2,108 | 136 | 2 | 2 |
232 | 3,866 | 283 | 2 | 3 |
232 | 5,036 | 312 | 3 | 1 |
232 | 32,428 | 2,229 | 3 | 2 |
232 | 102,531 | 7,411 | 3 | 3 |
@@ 3.2 Dataset Generation
An example of JSON’s PathCompose is shown in Fig. 2 to illustrate the dataset generation process. The generation process mainly entails constructing an abstract structure tree, manually drafting question templates, and developing corresponding answer discovery algorithms. The first step of the generation process is to define the complexity of the problem, characterized by depth, width, and column (Col), as well as its type, including task and language. During the construction of the abstract tree, depth represents the depth of the tree, width indicates the number of children for each non-leaf node, and Col specifies the number of fields associated with each node. When constructing the question template, predefined templates are retrieved based on the specified task. Finally, during sample generation, the selected task is used to identify the corresponding ground truth according to specific rules, and both the abstract tree and the ground truth are translated into the selected language.
Eight task categories have been delineated for eight languages, as detailed in Fig. 3(b). 29 rules and question templates have been formulated for these tasks, with the specific rule templates detailed in Sec. E in the Appendix. Each sample in the dataset comprises four main fields: “Reference”, “Question”, “Requirement” and “Answer”. We give examples for each language and task in Sec. D in the Appendix.
@@ 3.3 Statistic Information
StrucText-Eval has assembled two datasets. StrucText-Eval-Test comprises 3,712 samples, and StrucText-Eval-Test-Hard comprises 2,088 samples, each of the 29 specific tasks for eight languages as depicted in Fig. 3(a). Detailed statistics regarding the number of samples, lengths, and complexity levels across all tasks, languages, and difficulties are detailed in Tab. 1.
@@ 4.1 Models
We tested six Open-Source LLMs in both StrucText-Eval Test and Test-Hard Suite, and we use the short name (in the bracket) of these LLMs in the experiments: Qwen/Qwen2-7B-Instruct (Qwen-2-7B), Qwen/Qwen2-72B-Instruct (Qwen-2-72B), meta-llama/Meta-Llama-3.1-8B-Instruct (Llama-3.1-8B), meta-llama/Meta-Llama-3.1-72B-Instruct (Llama-3.1-70B), meta-llama/Meta-Llama-3.1-405B-Instruct (Llama-3.1-405B), mistralai/Mistral-7B-Instruct-v0.2 (Mistral-0.2-7B)
Considering the huge expense of using an API-based model, we only tested six Close-Source LLMs in StrucText-Eval-Hard: gpt-4o-2024-08-06 (gpt-4o), gpt-4o-mini-2024-07-18 (gpt-4o-mini), gemini-1.5-pro(gemini-1.5-pro), gemini-1.5-flash(gemini-1.5-flash), GLM-4-Plus (glm-4-plus), GLM-4-Flash (glm-4-flash).
@@ 4.2 Prompt-based Method
We also evaluated the impact of different prompt designs on the performance of LLMs by utilizing six distinct prompt configurations in the main experiments. Detailed implementation of these prompts can be found in Sec. C in the Appendix. The six primary prompt settings are as follows:
Naive: This configuration involves a straightforward input of “Context”, “Question”, and “Options” into the LLMs to generate responses. Self-Chain-of-Thought (Self-CoT) Kojima et al. (2022): This approach incorporates a step-by-step reasoning prompt to guide the model through logical reasoning. Plan-and-Solve CoT (PS-CoT) Wang et al. (2023): This method emphasizes problem decomposition before solving, encouraging the model to first break down the problem before generating a solution. With Hint (w/ hint): In this setting, manually curated hints are provided to the model to observe its performance when additional information is injected. Since this approach introduces supplementary data, it is delineated by a dashed line from other methods in Table 5. Few-Shot Demonstration: involves appending few training data directly to the prompt. The Simple Few-Shot Demonstration uses only the shortest examples from the training set as few-shot demonstrations.
@@ 4.3 Evaluation Method
We use the RougeL metric Lin (2004) to assess the degree of character-level similarity between model outputs in the main content of this paper. Sometimes, the task requires the LLM to generate the entire reasoning path leading to the answer, which results in high RougeL scores. So, we assign a score of 0 if the RougeL score falls below 0.75.
Additionally, we present the results of other evaluation metrics, including LLM-as-Judge-Score Zheng et al. (2023), BLEU Papineni et al. (2002), and Exact Match, in Tab. 4 in the Appendix. Furthermore, we conduct a consistency analysis across these metrics compared to human judgments, as shown in Fig. 6.
@@ 5 Analysis
Qwen2-7B | Base | 70.4 | 68.8 | 68.0 | 54.5 | 83.5 | 68.9 | 57.6 | 68.5 | 48.5 | 74.2 | 49.2 | 72.4 | 79.5 | 78.4 | 47.7 | 93.2 | 30.0 |
Self-CoT | 12.8 | 1.5 | 1.5 | 9.1 | 29.0 | 4.5 | 3.6 | 3.5 | 4.5 | 6.4 | 6.1 | 8.1 | 27.3 | 26.1 | 2.3 | 6.8 | 17.2 | |
PS-CoT | 31.7 | 31.7 | 19.4 | 20.1 | 67.0 | 36.4 | 25.8 | 24.9 | 9.8 | 19.8 | 32.6 | 34.1 | 63.6 | 60.2 | 25.0 | 72.7 | 29.1 | |